# MDNS: Masked Diffusion Neural Sampler via Stochastic Optimal Control

### v1

@@ 4.2 Potts model on Square Lattice
Similar to the training for the Ising model, we use $\mathcal{F}_{\mathrm{WDCE}}$ and train on each temperature for $100\mathrm{k}$ steps, among which $30\mathrm{k}$ steps are in the warm-up phase for training $\beta_{\mathrm{critical}}$ and $\beta_{\mathrm{low}}$. We report quantitative metrics (defined in ˜32 and 30) in Tab.˜3 and similarly plot the estimated average 2-point correlation (defined in (31) between a varying distance $r$ in Fig.˜2. More results can be found in App.˜E. Again, our approach recovers a distribution with a correlation function with decay patterns well aligned with the ground truth, while other baselines all fail to do so. It is worth noting that the MH algorithm cannot mix even after a continuous simulation of more than $20$ hours, while it remains a successful approach for all Ising model experiments. Such a difference likely originates from the exponential increase in the state space cardinality $|\mathcal{X}_{0}|$ despite the system size $L$ remaining the same. For example, a $4\times 4$ Ising model has around $60\mathrm{k}$ distinct states, while a $4\times 4$ Potts model with $q=3$ has $40\mathrm{M}$ distinct states, which clearly highlights the increased difficulty of Potts model sampling. Nevertheless, our approach manages to succeed with a moderate increase in training iterations. Moreover, the evaluation shows that MDNS has the best performance across almost all metrics and temperatures, suggesting our framework and learning objectives are highly scalable.
@@ 5 Conclusion, Limitations and Future Directions
This paper introduces Masked Diffusion Neural Sampler (MDNS), a novel framework for training discrete neural samplers based on stochastic optimal control and masked diffusion models. While MDNS has proven effective for distributions arising in statistical physics, its performance on other families of discrete distributions such as those on graphs or related to combinatorial optimization is unknown. Theoretically, we conjecture that our interpolation between the masked and the target distribution possesses superior properties compared to the geometric annealing $(\pi_{\eta}\propto\mathrm{e}^{-\eta U})_{\eta\in[0,1]}$ used in LEAPS [HAJ25]. A rigorous theoretical analysis of annealing paths in discrete spaces, potentially building upon insights in continuous spaces [GTC25a, GTC25], remains a key area for future investigation.
@@
WG and MT thank Peter Holderrieth and Michael Albergo for insightful discussions on the paper [HAJ25] during ICLR 2025. YZ and MT are grateful for partial support by NSF Grants DMS-1847802 and DMS-2513699, DOE Grant DE-NA0004261, Cullen-Peck Scholarship, and Emory-GT AI.Humanity Award. WG, JC, and YC acknowledge support from NSF Grants ECCS-1942523 and DMS-2206576.
@@ Appendix A Related Works
Recently, a growing trend of research directions has been focusing on training neural samplers [Noé+19, He+25], where one amortizes the sampling procedure through learning a neural network. Most of the works have focused on learning distributions on the continuous Euclidean space $\mathbb{R}^{d}$. One of the main approaches is to learn a trainable drift term that drives an SDE to approximate the time-reversal of a pre-selected noising process that converts any target distribution $\pi$ to a noise distribution. This includes methods like Path Integral Sampler (PIS, [ZC22]), Neural Schrödinger-Föllmer Flows (NSFS, [Var+23]), Denoising Diffusion Samplers (DDS, [VGD23]), time-reversed DIffusion Sampler (DIS, [BRU24]), iterated Denoising Energy Matching (iDEM, [AS+24]), Diffusion-PINN Sampler [Shi+24a], Adjoint Sampling [Hav+25, Liu+25a, Cho+25], etc. Another popular approach trains the process to transport from a prior distribution $\pi_{0}$ to $\pi$ along a sequence of marginal distributions whose unnormalized densities are all available, such as the geometric interpolation $\pi_{t}\propto\pi_{0}^{1-t}\pi_{1}^{t}$. Methods of this type include the paper [MF23], Controlled Monte Carlo Diffusion (CMCD, [Var+24]), Liouville Flow Importance Sampler (LFIS [TPL24]), Non-Equilibrium Transport Sampler (NETS, [AVE25]), Sequential Controlled Langevin Diffusions (SCLD, [Che+25b]), the paper [Che+25], underdamped diffusion bridges [Ble+25], Accelerated Parallel Tempering (APT, [Zha+25]), Progessive Tempering with Diffusion (PTSD, [Ris+25]), etc. There are also works motivated by annealed importance sampling [Nea01] and Jarzynski equality [Jar97, VJ08], including Monte Carlo Diffusion (MCD, [Dou+22]) and Langevin Diffusion Variational Inference (LDVI, [GD23]). Despite the progress, these methods are restricted to the setting of $\mathbb{R}^{d}$ and cannot be easily extended to sampling discrete distributions. This distinguishes our proposed MDNS from all the aforementioned approaches.
@@
Diffusion models [SD+15, HJA20, SME21, Son+21] have been achieving state-of-the-art performances on generative modeling of various data modalities [Wat+23, Zhu+25b, Ess+24, Jin+25, Zhu+25, Chi+23, HRT24, Roj+25a, Che+25a]. As a natural generalization of diffusion models to discrete state space, discrete diffusion models [Aus+21, Cam+22, CZH23, Sun+23a, LME24, Gat+24, Sha+25] have gradually attracted the attention of the research community as a competent method for generative modeling of general discrete sequence data. Discrete diffusion models have seen successful applications in numerous tasks, including image synthesis [Cha+22, Bai+25, Shi+25], text generation [Nie+25, Gon+25, Zhe+24, Arr+25], protein [Gru+23, Ala+23, Hay+25, Wan+24, Che+25c, TZC25], graph [Xu+24, Li+25], etc. Motivated by the huge empirical success of discrete diffusion models in practice, researchers have also worked on the theory of discrete diffusion to understand the key behind its success [Ben+24, CY25, Ren+25, Fen+25, Kim+25, RRY25, HRT25]. An extensive portion of the literature is also devoted to the development of techniques to enhance further the effectiveness of discrete diffusion models, such as guidance [Sch+25, Guo+24, Nis+25, Roj+25], distillation [Zhu+25a, DG25], training [Zha+25a, Liu+25], planning [Par+25, Liu+25b], and inference-time scaling [Wan+25a].
@@ C.3 Omitted Proofs in the Main Text
Therefore, we have $\mathbb{P}^{0}_{T|0}(n|\mathbf{M})=\frac{1}{N}\left(1-\exp\left(\int_{0}^{T}\gamma(r)\mathrm{d}r\right)\right)$ for $n\in\{1,...,N\}$. Since the initial sample is completely masked, to guarantee $\mathbb{P}^{0}_{T}=p_{\mathrm{unif}}$, it suffices to make $\int_{0}^{T}\gamma(r)\mathrm{d}r=\infty$. Moreover, as the initial distribution is a point mass, it is obvious that $X_{0}$ and $X_{T}$ are independent, i.e., $\mathbb{P}^{0}$ is memoryless. ∎

### v2

@@ 4.2 Potts model on Square Lattice
Similar to the training for the Ising model, we use $\mathcal{F}_{\mathrm{WDCE}}$ and train on each temperature for $100\mathrm{k}$ steps, among which $30\mathrm{k}$ steps are in the warm-up phase for training $\beta_{\mathrm{critical}}$ and $\beta_{\mathrm{low}}$. We report quantitative metrics (defined in ˜32 and 30) in Tab.˜3 and similarly plot the estimated average 2-point correlation (defined in (31) between a varying distance $r$ in Fig.˜2. More results can be found in App.˜E. Again, our approach recovers a distribution with a correlation function whose decay patterns are well aligned with the ground truth, whereas other baselines fail to do so. It is worth noting that the MH algorithm cannot mix even after a continuous simulation of more than $20$ hours, while it remains a successful approach for all Ising model experiments. Such a difference likely originates from the exponential increase in the state space cardinality $|\mathcal{X}_{0}|$ despite the system size $L$ remaining the same. For example, a $4\times 4$ Ising model has around $60\mathrm{k}$ distinct states, while a $4\times 4$ Potts model with $q=3$ has $40\mathrm{M}$ distinct states, which clearly highlights the increased difficulty of Potts model sampling. Nevertheless, our approach succeeds with a moderate increase in training iterations. Moreover, the evaluation shows that MDNS performs best across almost all metrics and temperatures, suggesting that our framework and learning objectives are highly scalable.
@@ 5 Conclusion, Limitations and Future Directions
This paper introduces Masked Diffusion Neural Sampler (MDNS), a novel framework for training discrete neural samplers based on stochastic optimal control and masked diffusion models. While MDNS has proven effective for distributions arising in statistical physics, its performance on other families of discrete distributions, such as those on graphs or related to combinatorial optimization, is unknown. We also propose that the framework can be extended to fine-tune pretrained discrete diffusion models given a (possibly non-differentiable) reward function, which is left for future work [Zhu+25c]. Theoretically, we conjecture that our interpolation between the masked and the target distribution possesses superior properties compared to the geometric annealing $(\pi_{\eta}\propto\mathrm{e}^{-\eta U})_{\eta\in[0,1]}$ used in LEAPS [HAJ25]. A rigorous theoretical analysis of annealing paths in discrete spaces, potentially building upon insights in continuous spaces [GTC25a, GTC25], remains a key area for future investigation.
@@
WG and MT thank Peter Holderrieth and Michael Albergo for insightful discussions on the paper [HAJ25] during ICLR 2025. YZ and MT are grateful for partial support by NSF Grants DMS-1847802, DMS-2513699, DOE Grants NA0004261, SC0026274, and Richard Duke Fellowship. WG, JC, and YC acknowledge support from NSF Grants ECCS-1942523, DMS-2206576, and CMMI-2450378.
@@ Appendix A Related Works
Recently, a growing trend of research directions has been focusing on training neural samplers [Noé+19, He+25], where one amortizes the sampling procedure through learning a neural network. Most of the work has focused on learning distributions over the continuous Euclidean space $\mathbb{R}^{d}$. One of the main approaches is to learn a trainable drift term that drives an SDE to approximate the time-reversal of a pre-selected noising process that converts any target distribution $\pi$ to a noise distribution. This includes methods like Path Integral Sampler (PIS, [ZC22]), Neural Schrödinger-Föllmer Flows (NSFS, [Var+23]), Denoising Diffusion Samplers (DDS, [VGD23]), time-reversed DIffusion Sampler (DIS, [BRU24]), iterated Denoising Energy Matching (iDEM, [AS+24]), Diffusion-PINN Sampler [Shi+24a], Adjoint Sampling [Hav+25, Liu+25a, Cho+25], etc. Another popular approach trains the process to transport from a prior distribution $\pi_{0}$ to $\pi$ along a sequence of marginal distributions whose unnormalized densities are all available, such as the geometric interpolation $\pi_{t}\propto\pi_{0}^{1-t}\pi_{1}^{t}$. Methods of this type include the paper [MF23], Controlled Monte Carlo Diffusion (CMCD, [Var+24]), Liouville Flow Importance Sampler (LFIS [TPL24]), Non-Equilibrium Transport Sampler (NETS, [AVE25]), Sequential Controlled Langevin Diffusions (SCLD, [Che+25b]), the paper [Che+25], underdamped diffusion bridges [Ble+25], Accelerated Parallel Tempering (APT, [Zha+25]), Progessive Tempering with Diffusion (PTSD, [Ris+25]), etc. There are also works motivated by annealed importance sampling [Nea01] and Jarzynski equality [Jar97, VJ08], including Monte Carlo Diffusion (MCD, [Dou+22]) and Langevin Diffusion Variational Inference (LDVI, [GD23]). Despite the progress, these methods are restricted to the setting of $\mathbb{R}^{d}$ and cannot be easily extended to sampling discrete distributions. This distinguishes our proposed MDNS from all the aforementioned approaches.
@@
Diffusion models [SD+15, HJA20, SME21, Son+21] have been achieving state-of-the-art performances on generative modeling of various data modalities [Wat+23, Zhu+25b, Ess+24, Jin+25, Zhu+25, Chi+23, HRT24, Roj+25a, Che+25a, Ren+25b]. As a natural generalization of diffusion models to discrete state space, discrete diffusion models [Aus+21, Cam+22, CZH23, Sun+23a, LME24, Gat+24, Sha+25] have gradually attracted the attention of the research community as a competent method for generative modeling of general discrete sequence data. Discrete diffusion models have seen successful applications in numerous tasks, including image synthesis [Cha+22, Bai+25, Shi+25], text generation [Nie+25, Gon+25, Zhe+24, Arr+25], protein [Gru+23, Ala+23, Hay+25, Wan+24, Che+25c, TZC25], graph [Xu+24, Li+25], combinatorial optimization [SHL24, San+25] etc. Motivated by the huge empirical success of discrete diffusion models in practice, researchers have also worked on the theory of discrete diffusion to understand the key behind its success [Ben+24, CY25, Ren+25, Fen+25, Kim+25, RRY25, HRT25]. An extensive portion of the literature is also devoted to the development of techniques to enhance further the effectiveness of discrete diffusion models, such as guidance [Sch+25, Guo+24, Nis+25, Roj+25], distillation [Zhu+25a, DG25], training [Zha+25a, Liu+25], planning [Par+25, Liu+25b], and inference-time scaling [Wan+25a, Tan+25].
@@ C.3 Omitted Proofs in the Main Text
Therefore, we have $\mathbb{P}^{0}_{T|0}(n|\mathbf{M})=\frac{1}{N}\left(1-\exp\left(\int_{0}^{T}\gamma(r)\mathrm{d}r\right)\right)$ for $n\in\{1,...,N\}$. Since the initial sample is completely masked, $\int_{0}^{T}\gamma(r)\mathrm{d}r=\infty$ guarantees $\mathbb{P}^{0}_{T}=p_{\mathrm{unif}}$. Moreover, as the initial distribution is a point mass, it is obvious that $X_{0}$ and $X_{T}$ are independent, i.e., $\mathbb{P}^{0}$ is memoryless. ∎
@@ F.3 Experiments
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We believe the main claims made in the paper accurately reflect our contributions.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We have discussed the limitations in Sec.˜5.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: All theoretical results have a complete proof in the appendix. We have checked the mathematical derivation and affirmed that they are correct.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: All information regarding of reproducing the experiments can be found in the appendix.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Our experiments do not rely on any data. The code repository can be found in the abstract.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: All experimental details can be found in the appendix.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: We do not report error bars in the paper.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: The information on the computing resources for the experiments can be found in the appendix.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Our research completely complies with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: The discussion of broader impacts can be found after the references.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [NA]
  54. Justification: We believe our paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We have mentioned the original papers that produced the code in the appendix.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [NA]
  64. Justification: Our paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [NA]
  69. Justification: Our paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [NA]
  74. Justification: Our paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [NA]
  79. Justification: The the core method development in our paper does not involve LLMs as any important, original, or non-standard components.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.